[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Oct 18, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Oct 19 04:20:53 UTC 2016


HSU Shift Log for Wed Oct 19 2016 00:20:51 GMT-0400 (EDT)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: Issue with submitting guidance at 00z. System hung after  
submitting it for all invests. Called SDM and found out that  
supercomputer "tide" is down for maintenance. Monica indicated that our  
ability to ssh into the supercomputer is not working. She will contact SDM.  
Until this is fixed, after submitting guidance and ATCF hangs (terminal  
window asking for password) hit Control C twice and it will get the ATCF  
past that portion of the script. This should allow guidance to be  
submitted. No problem with the retrieval process. This may need to be done  
at 12z until the issuance is resolve. These instructions have been left on  
the ATCF.

N-AWIPS/AWIPS2 issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161019/b733954d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list